A rectangular billboard 6 feet in height stands in a field so that its bottom is 13 feet above the ground. A nearsighted cow with eye level at 4 feet above the ground stands x

x

 feet from the billboard. Express θ

θ

, the vertical angle subtended by the billboard at her eye, in terms of x

x

. Then find the distance x


x


 the cow must stand from the billboard to maximize θ

θ

.



1
Expert's answer
2022-01-10T16:19:27-0500
"\theta(x)=\tan^{-1}\dfrac{13+6-4}{x}-\tan^{-1}\dfrac{13-4}{x}"

"\theta(x)=\tan^{-1}\dfrac{15}{x}-\tan^{-1}\dfrac{9}{x}, x>0"

Take the derivative with respect to "x"


"\theta '(x)=\dfrac{1}{1+(\dfrac{15}{x})^2}(-\dfrac{15}{x^2})-\dfrac{1}{1+(\dfrac{9}{x})^2}(-\dfrac{9}{x^2})"

"=\dfrac{9}{x^2+81}-\dfrac{15}{x^2+225}"

Find the critical number(s)


"\theta '(x)=0=>\dfrac{9}{x^2+81}-\dfrac{15}{x^2+225}=0"

"\dfrac{9}{x^2+81}=\dfrac{15}{x^2+225}"

"3(x^2+225)=5(x^2+81)"

"2x^2=270"

"x^2=135"

"x=\pm\sqrt{135}"

"x=\pm3\sqrt{15}"

Since "x>0" we take "x=3\sqrt{15}."

If "0<x<3\sqrt{15}, \theta'(x)>0, \theta(x)" increases.

If "x>3\sqrt{15}, \theta'(x)<0, \theta(x)" decreases.

The cow must stand at "3\sqrt{15}" feet from the billboard to maximize "\u03b8."